OpenStudy (anonymous):

I would like some help calculating the flux of neutrinos through the Earth. I am given the energy released in neutrinos, 10E53 erg, and the mean energy 10MeV per neutrino. I am asked how many pass through a cm^2. I know... • Flux = N/(At) = E/(At) * N/E • 10E53 erg = 6.24 x 10^64 eV, which leads to... • 6.24x10^57 neutrinos. I'm feeling lost on the next step as I believe I need the energy that passes through some area on Earth, but I'm not sure. It seems to simple and wrong to simply divide # of neutrinos by (1 x 10^-4 m)*s Thanks for your help.

OpenStudy (anonymous):

Can you explain why this is your solution?

OpenStudy (anonymous):

The energy released in neutrinos/per second per unit area is coming from the Sun. At earth's distance d= 1AU = 1.5 * 10^10 cm, the flux would be this energy divided by the area of the sphere of radius d. hence \[energy flux = E/(4*\pi*d^2)\] so neutrino flux = energy flux / energy per neutrino

OpenStudy (anonymous):

Right, but the question is asking for the flux through an area of one square cm. Isn't that flux you give the flux through a sphere around the Sun, not the flux through some area of the Earth?

OpenStudy (anonymous):

If my description above is lacking, maybe seeing the original problem would help. Please refer to problem (4) on this page: http://www.physics.rutgers.edu/~pryor/ph442/hw1.pdf Thanks for your help.

OpenStudy (anonymous):

mate the d that I wrote above in the equation, is the distance of earth from the sun. now imagine, earth is just a point mass in the big sphere thats formed with sun as the center and d as the radius. The total neutrino energy coming out from the sun would be passing through this sphere. hence, you have to divide this energy from the area of this sphere to get the energy per unit area.